# February 12, 2024

### v1

@@ VI.4 Convergence Metrics
Sampler HT in Loop Generation Evaluation Density Estimator Overhead (sec) (sec) Training (sec) (sec) CMAES False $0.48\pm 0.16$ $0.43\pm 0.15$ $(3.9\pm 23)\times 10^{-3}$ $0.055\pm 0.043$ True $1.6\pm 0.3$ $1.6\pm 0.2$ $(1.9\pm 22)\times 10^{-3}$ $0.068\pm 0.048$ Random False $0.33\pm 0.06$ $0.31\pm 0.06$ N/A $0.015\pm 0.003$ True $1.4\pm 0.2$ $1.3\pm 0.2$ N/A $0.019\pm 0.012$
Table 4: Comparison of times taken by different parts of the loop for CMAES and random sampler with and without HiggsTools in the loop. The times refer to a generation of 12 points in every case. The values presented are the mean $\pm$ one standard deviation over the scans falling in each of the four categories.
@@ A.4 Unitarity
A valid point in the parameter space must also satisfy the perturbative unitarity constraints. This can be expressed in terms of constraints on the $\lambda_{i}$ potential parameters. For the different symmetry constrained 3HDM these are fully given in [11] to which we refer the reader for further details.

### v2

@@ VI.4 Convergence Metrics
Sampler HiggsTools in Loop Generation Evaluation Density Estimator Overhead CMAES False $0.48\pm 0.16$ $0.43\pm 0.15$ $(3.9\pm 23)\times 10^{-3}$ $0.055\pm 0.043$ True $1.6\pm 0.3$ $1.6\pm 0.2$ $(1.9\pm 22)\times 10^{-3}$ $0.068\pm 0.048$ Random False $0.33\pm 0.06$ $0.31\pm 0.06$ N/A $0.015\pm 0.003$ True $1.4\pm 0.2$ $1.3\pm 0.2$ N/A $0.019\pm 0.012$
Table 4: Comparison of times, in seconds, taken by different parts of the loop for CMAES and random sampler with and without HiggsTools in the loop. The times refer to a generation of 12 points in every case. The values presented are the mean $\pm$ one standard deviation over the scans falling in each of the four categories.
@@ A.4 Unitarity
A valid point in the parameter space must also satisfy the perturbative unitarity constraints. These can be expressed in terms of constraints on the $\lambda_{i}$ parameters of the potential. For this end we require that the eigenvalues of all the two-by-two scattering matrices are bounded by partial wave unitarity. In [11] an algorithm was proposed to efficiently obtain the scattering matrices. For the different symmetry constrained 3HDM these are fully given in [12] to which we refer the reader for further details.